Can you please help me solve using the substitution method?
x + 7y = 16
-3x + 6y = 33

Answer by checkley71(8403)

X+7Y=16 THEN
X=16-7Y NOW SUBSTITUTE (16-7Y) FOR X IN THE OTHER EQUATION THUS
-3(16-7Y)+6Y=33
-48+21+6Y=33
6Y=33+48-21
6Y=60
Y=60/6
Y=10 SOLUTION
X=16-7*10
X=16-70
X=-54
PROOF
-54+7*10=16
-54+70=16
16=16